# VFM-VLM: Vision Foundation Model and Vision Language Model based Visual Comparison for 3D Pose Estimation

### v1

@@ Abstract
Vision Foundation Models (VFMs) and Vision Language Models (VLMs) have revolutionized computer vision by providing rich semantic and geometric representations. This paper presents a comprehensive visual comparison between CLIP based and DINOv2 based approaches for 3D pose estimation in hand object grasping scenarios. We evaluate both models on the task of 6D object pose estimation and demonstrate their complementary strengths: CLIP excels in semantic understanding through language grounding, while DINOv2 provides superior dense geometric features. Through extensive experiments on benchmark datasets, we show that CLIP based methods achieve better semantic consistency, while DINOv2 based approaches demonstrate competitive performance with enhanced geometric precision. Our analysis provides insights for selecting appropriate vision models for robotic manipulation and grasping, picking applications.
@@ I Introduction
  • Insights into the trade offs between semantic understanding and geometric precision
@@ II Related Work
II-A Traditional 3D Pose Estimation
@@
II-B Deep Learning for Pose Estimation
@@
II-C Vision Foundation Models
@@ III-A Problem Formulation
III-B CLIP Based Architecture
The CLIP based approach leverages semantic understanding through language grounding:
Feature Extraction: The CLIP vision encoder processes the input image to extract multi scale features: $\mathbf{f}_{clip}=\text{CLIP}_{\text{vis}}(I)$
Semantic Enhancement: Text prompts describing object affordances and grasping intentions are encoded: $\mathbf{t}_{sem}=\text{CLIP}_{\text{text}}(\text{"A hand grasping a driller"})$
Cross Modal Fusion: Visual and language features are fused through attention mechanisms to produce semantically aware representations.[10]
@@ III-B CLIP Based Architecture
III-C DINOv2 Based Architecture
The DINOv2 based approach emphasizes dense geometric features:
Dense Feature Extraction: DINOv2 produces high resolution feature maps with strong spatial correspondence: $\mathbf{f}_{dino}=\text{DINOv2}(I)$
@@ IV-B Quantitative Comparison
IV-C Visual Analysis
@@
DINOv2 Based Results (Figures 2 and 3): The second figure shows DINOv2’s predictions with multiple 3D bounding boxes overlaid on the scene. The dense geometric features enable accurate localization of multiple objects simultaneously, with tighter alignment to ground truth poses. The superior ADD distance (28.45mm) reflects DINOv2’s strength in geometric reasoning.
Figure 3 provides a complementary 3D visualization of the DINOv2 backbone’s performance, showing the complete scene reconstruction with RGB point cloud data. The predicted 3D bounding boxes (displayed in red and green) demonstrate precise spatial localization with accurate depth estimation. The coordinate system spans approximately 1800mm in height (Z-axis) and ±600mm in the XY plane, showcasing the model’s capability to handle real world metric measurements. This visualization emphasizes DINOv2’s strength in dense geometric understanding, where each point in the cloud maintains spatial consistency with the predicted object poses.
IV-D Qualitative Observations
Semantic vs. Geometric Trade off: CLIP excels at understanding object affordances and grasp intentions but sacrifices geometric precision. DINOv2 provides superior spatial accuracy but lacks explicit semantic reasoning. As evidenced in Figure 3, DINOv2’s dense point cloud representation enables millimeter level precision in 3D space.
@@ V-A Complementary Strengths
Our experiments reveal that CLIP and DINOv2 address different aspects of the pose estimation problem. CLIP’s language grounding enables task conditioned reasoning: ”grasp the driller by the handle” produces different pose hypotheses than grasp the driller by the body. DINOv2’s dense correspondences provide the geometric precision necessary for accurate pose recovery.

### v2

@@ Abstract
Vision Foundation Models (VFMs) and Vision Language Models (VLMs) have revolutionized computer vision by providing rich semantic and geometric representations. This paper presents a comprehensive visual comparison between CLIP based and DINOv2 based approaches for 3D pose estimation in hand object grasping scenarios. We evaluate both models on the task of 6D object pose estimation and demonstrate their complementary strengths: CLIP excels in semantic understanding through language grounding, while DINOv2 provides superior dense geometric features. Through extensive experiments on benchmark datasets, we show that CLIP based methods achieve better semantic consistency, while DINOv2 based approaches demonstrate competitive performance with enhanced geometric precision. Our analysis provides insights for selecting appropriate vision models for robotic manipulation and grasping, picking applications.[1]
@@ I Introduction
  • Insights between semantic understanding and geometric precision
@@ II Related Work
II-A Taxonomy of Pose Estimation
3D pose estimation can be systematically categorized along multiple dimensions, forming a comprehensive taxonomy that guides our understanding of the field’s evolution.
Input Modality: Methods are divided into RGB-only approaches, which leverage color information and learned representations, RGB-D methods that exploit depth sensors for geometric constraints, and multi-modal approaches combining various sensor inputs. Recent VFM based methods primarily operate on RGB input but extract rich geometric and semantic features.
Output Representation: Direct pose regression methods predict 6D pose parameters (rotation and translation)[5] end-to-end. Keypoint-based approaches first detect 2D projections of 3D keypoints, then solve PnP for pose recovery. Dense correspondence methods[6] establish pixel-to-model mappings before pose optimization. Hybrid methods combine multiple representations for robustness.
Learning Paradigm: Supervised methods require extensive labeled pose annotations. Self-supervised approaches leverage geometric constraints (epipolar geometry, photometric consistency)[7] or reconstruction losses. Foundation models introduce a new paradigm of pre-training on massive unlabeled datasets followed by task specific fine-tuning, bridging the gap between pure supervision and self-supervision.
Architecture: CNN based methods (PoseCNN, DenseFusion)[8] [9] dominated early deep learning approaches. Transformer-based architectures capture global context and long range dependencies. Point cloud networks (PointNet, PointNet++)[10] process 3D geometric data directly. VFM based approaches leverage pre-trained vision transformers with billions of parameters, representing the current frontier.
This taxonomy reveals that VFM aided methods represent a convergence point: they primarily use RGB input, produce multiple output representations (keypoints, dense features, direct pose), leverage self-supervised pre-training at scale, and employ transformer architectures. Our work positions CLIP and DINOv2 within this taxonomy as representatives of language grounded and pure visual self-supervised paradigms, respectively.
II-B Traditional 3D Pose Estimation
@@
II-C Deep Learning for Pose Estimation
@@
II-D Vision Foundation Models
@@ III-A Problem Formulation
III-B Model Architectures
Figure 1: CLIP Architecture: Dual-encoder framework with Vision Transformer (ViT-B/32) for image encoding and Text Transformer for language encoding. Both encoders project to a shared 512-dimensional embedding space where contrastive learning aligns matched image-text pairs.
Figure 2: DINOv2 Architecture: Self-supervised Vision Transformer (ViT-B/14) using student-teacher framework with self-distillation. The model produces dense patch-level features (768-dim) with strong spatial correspondence, ideal for geometric reasoning tasks.
CLIP (Figure 1) employs a dual-encoder architecture trained through contrastive learning on 400M image-text pairs, producing semantically-aligned visual and language representations. DINOv2 (Figure 2) leverages self-supervised learning on 142M images through self-distillation, generating dense geometric features without language supervision. The key distinction lies in their training paradigms: CLIP learns cross-modal semantic alignment while DINOv2 focuses on intra-modal spatial correspondence.
III-C CLIP-Based Architecture for 6D Pose
The CLIP-based approach leverages semantic understanding through language grounding:
Feature Extraction: The CLIP vision encoder processes the input image to extract multi-scale features: $\mathbf{f}_{clip}=\text{CLIP}_{\text{vis}}(I)$
Cross-Modal Fusion: Visual and language features are fused through attention mechanisms to produce semantically-aware[14] representations.
@@ III-C CLIP-Based Architecture for 6D Pose
III-D DINOv2-Based Architecture for 6D Pose
The DINOv2-based approach emphasizes dense geometric features:
Dense Feature Extraction: DINOv2 produces high-resolution feature maps with strong spatial[15] correspondence: $\mathbf{f}_{dino}=\text{DINOv2}(I)$
@@ IV-B Quantitative Comparison
IV-C Qualitative Observations & Visual Analysis
@@
DINOv2 Based Results (Figures 4 and 5): The second figure shows DINOv2’s predictions with multiple 3D bounding boxes overlaid on the scene. The dense geometric features enable accurate localization of multiple objects simultaneously, with tighter alignment to ground truth poses. The superior ADD distance (28.45mm) reflects DINOv2’s strength in geometric reasoning.
Figure 5 provides a complementary 3D visualization of the DINOv2 backbone’s performance, showing the complete scene reconstruction with RGB point cloud data. The predicted 3D bounding boxes (displayed in red and green) demonstrate precise spatial localization[17] with accurate depth estimation. The coordinate system spans approximately 1800mm in height (Z-axis) and ±600mm in the XY plane, showcasing the model’s capability to handle real world metric measurements. This visualization emphasizes DINOv2’s strength in dense geometric understanding, where each point in the cloud maintains spatial consistency with the predicted object poses.
Semantic vs. Geometric: CLIP excels at understanding object affordances and grasp intentions but sacrifices geometric precision. DINOv2 provides superior spatial accuracy but lacks explicit semantic reasoning. As evidenced in Figure 5, DINOv2’s dense point cloud representation enables millimeter level precision in 3D space.
@@ V-A Complementary Strengths
Our experiments reveal that CLIP and DINOv2 address different aspects of the pose estimation problem. CLIP’s language grounding enables task conditioned reasoning: ”grasp the driller by the handle”[18] produces different pose hypotheses than grasp the driller by the body. DINOv2’s dense correspondences provide the geometric precision[19] necessary for accurate pose recovery.